[Hallc_running] [New Logentry] Hall C RC Update (Sept 10)

brads at jlab.org brads at jlab.org
Thu Sep 10 12:05:01 EDT 2020

Logentry Text:
<p><a class="new" title="RC Update and Run plan for Sep 10,2020" href="https://hallcweb.jlab.org/wiki/index.php?title=RC_Update_and_Run_plan_for_Sep_10,2020&action=edit&redlink=1">RC Update and Run plan for Sep 10,2020</a></p><h1><span id="General_updates" class="mw-headline">General updates</span></h1><p>(See <a title="D2n Run Plan 2020" href="https://hallcweb.jlab.org/wiki/index.php/D2n_Run_Plan_2020">d2n Run Plan 2020</a> for broader status/overview.)</p><ul><li>Production running on Transverse target; Kin-C, Kin-Z continues.<ul><li>General plan when this is done is to complete Kin-A, Kin-X (partially done already); then move on to Kin-B/Y and/or 1-pass running as time allows.</li></ul></li><li>Target Tommy polarization @ 270 is lower than we would like<ul><li>Maximum seen 37.5% <a class="external autonumber" href="https://logbooks.jlab.org/entry/3844366" rel="nofollow">[1]</a>; presently around 36%</li><li>QWP optimization restarted Wed swing; ongoing</li><li>For!
  comparison, maximum seen with target at 90 deg: 44% (9/5/2020 22:56:38)</li></ul></li><li>Extremely trippy beam last night, up to 35--40 trips/hour<ul><li>MCC took beam away at 9am to try to address BLM issues in the North, see if they can recover some cavities, and address issues with the 5th pass separator. Nominal 2 hour down, still down as of noon.</li></ul></li><li>Opportunistic Target Studies<ul><li>AFP loss study completed <a class="external autonumber" href="https://logbooks.jlab.org/entry/3845255" rel="nofollow">[2]</a></li><li>Temperature study @ 270deg in progress</li></ul></li></ul><p> </p><h2><span id="Notes_for_the_Future" class="mw-headline">Notes for the Future</span></h2><ul><li><strong>Moller</strong> measurement for <strong>Thursday DAY</strong> shift</li><li>Beam Studies planned for Tuesday, Sept 15<ul><li>Simona wants access at this time for SOLID cherenkov related work</li><li>Target work (install spectrum analyzer collimator?)</li></ul></li></ul!
 ><ul><li><a title="D2n Run Plan 2020" href="https://hallcweb.j!
 lab.org/wiki/index.php/D2n_Run_Plan_2020#1-pass_Running">1-pass running</a> tentatively scheduled for Sept 17/18</li></ul><h2><span id="Pending_Issues" class="mw-headline">Pending Issues</span></h2><ul><li>Target polarization for Tommy @ 270 is low (mid to high 30s)<ul><li>Spectrum analyzer collimator was broken at some point and present optics setup may be allowing transient light to interfere with spectrum measurement and QWP scans</li></ul></li><li>Beam energy drifts<ul><li>Hall C 5-pass polarization goes down as energy (HALLC:p) goes up <a class="external autonumber" href="https://logbooks.jlab.org/entry/3844493" rel="nofollow">[3]</a></li></ul></li><li>Trip rates<ul><li>Hall A Limited to 140 uA to Reduce ILM0L02 Trips <a class="external autonumber" href="https://logbooks.jlab.org/entry/3844844" rel="nofollow">[4]</a></li></ul></li><li>SHMS dipole LN2 loop oscillating <a class="external autonumber" href="https://logbooks.jlab.org/entry/3844119" rel="nofollow">[5]</a></l!
 i></ul><ul><li>pNMR<ul><li>Software / fitting issues? <a class="external autonumber" href="https://logbooks.jlab.org/entry/3844891" rel="nofollow">[6]</a></li><li>Summary for Tommy at 270 <a class="external autonumber" href="https://logbooks.jlab.org/entry/3844699" rel="nofollow">[7]</a>, at 90 <a class="external autonumber" href="https://logbooks.jlab.org/entry/3844044" rel="nofollow">[8]</a></li></ul></li></ul><h2><span id="Opportunistic_Jobs.2FAccess" class="mw-headline">Opportunistic Jobs/Access</span></h2><ul><li>Beam Studies planned for Tuesday, Sept 15<dl><dt>Target calibrations / tests</dt></dl><ul><li>Just before Moller measurement, and/or other known downtimes</li><li>See <a title="D2n Run Plan 2020" href="https://hallcweb.jlab.org/wiki/index.php/D2n_Run_Plan_2020#Target_Calibrations_and_Tests">D2n_Run_Plan_2020#Target_Calibrations_and_Tests</a></li></ul></li></ul><ul><li><dl><dt>SOLID Cherenkov Test -- Access request [Simona]</dt></dl><ul><li>She would prefer to !
 have access as soon as possible starting this Friday (Sep 11, Day Shift!
 ), to change the gas in the Cherenkov Tank during an opportunistic controlled access to the Hall during unexpected down times. NOTE: She may not need tech support, if she can use the ladder to reach fill the tank. \</li><li>Need ~1 hr without tech support. With tech support, may be 1-2 hrs.</li><li>Preferably, this work can be done during Day Shift, but it can also be done on Swing/Owl.)</li></ul></li></ul><h1><span id="Current_Run_Plan" class="mw-headline">Current Run Plan</span></h1><h2><span id="Wednesday_DAY_.E2.86.92_Thursday_OWL" class="mw-headline">Wednesday DAY → Thursday OWL</span></h2><ul><li>Continue production data (HMS Kin-C/SHMS Kin-Z) with pol 3He target, <strong>Transverse/270</strong> polarization with NMR every 5 hrs.</li></ul><h2><span id="Thursday_DAY_.E2.86.92_Thursday_SWING" class="mw-headline">Thursday DAY → Thursday SWING</span></h2><ul><li><strong>Moller measurement</strong> planned for Thursday DAY shift to begin around 9:30am</li><li>Before ha!
 nding over to Moller, the target experts will<ul><li>Take NMR/pNMR</li><li>Perform one of the opportunistic Target Study tests (AFP loss test [B4]?)</li></ul></li><li>Wait for Moller to complete and get production beam re-established (late-DAY, early-SWING)</li></ul><ul><li><strong>Continue</strong> production data (HMS Kin-C/SHMS Kin-Z) with pol 3He target, <strong>Transverse/270</strong> polarization with NMR every 5 hrs.</li></ul>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3845317
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200910/6403147e/attachment.html>

More information about the Hallc_running mailing list